A water sprinkler sprays water over a distance of 30 feet while rotating through an angle of 135 degrees. What area of the lawn recieves water?
Answer by stanbon(75887)

A water sprinkler sprays water over a distance of 30 feet while rotating through an angle of 135 degrees. What area of the lawn recieves water?
----
Assume the lawn is a circle with radius = 30 ft.
---
Area of the whole circle = (pi)*30^2 sq.ft.
----
Proportion:
x/900(pi) = 135/360
---
x = [900pi][135/360] = 337.5(pi) sq ft (area that receives water)
================================================
Cheers,
Stan H.